OpenStudy (anonymous):

give the ordered pair that satisfies the system of equations below... x= -5y-7 4x + 5y = 17 someone please help me! i really dont understand it. can you please indicate me what to do!

myininaya (myininaya):

ok you see x=-5y-7 plug that x into the other equation 4(-5y-7)+5y=17

myininaya (myininaya):

now solve for y

OpenStudy (anonymous):

first you would distribute 4 right?? omg this is highschool stuff and im only in middle school! -_-

myininaya (myininaya):

yes

myininaya (myininaya):

-20y-28+5y=17 now what

OpenStudy (anonymous):

plus 28 on both sides. meaning you wud cancel out -28 and add 28 to 17

OpenStudy (anonymous):

wait i got -25y - 28 = 17

OpenStudy (anonymous):

i combined like terms.

OpenStudy (anonymous):

then i ended up with -25y = 45

myininaya (myininaya):

-15y=45 y=-3

OpenStudy (anonymous):

ahhh thanks. my bad.

OpenStudy (anonymous):

then what do i do? i have to give the ordered pairs.

myininaya (myininaya):

what is x if y=-3

myininaya (myininaya):

remember x=-5y-7 and we know y=-3 so x=?

OpenStudy (anonymous):

yeah. so x = 8 ! omggg

OpenStudy (anonymous):

you've made my life so much easier

OpenStudy (anonymous):

my teacher taught it to me in an impossible way. thank you so much.

OpenStudy (anonymous):

you taught me in a different way.

myininaya (myininaya):

the ordered pair (x,y)=(8,-3)

OpenStudy (anonymous):

yeah. i got it! thank you so much!

myininaya (myininaya):

there are other ways to solve this types i like substitution the most thats what we did above

OpenStudy (anonymous):

my teacher told me to put it in y=mx + b form. but i didnt know how to do that.

myininaya (myininaya):

solve both of them for y

OpenStudy (anonymous):

what do you mean?

myininaya (myininaya):

x=-5y-7 add 7 to both sides x+7=-5y now divide by -5 on both sides (x+7)/(-5)=y y=x/(-5)+7/(-5) y=-x/5-7/5 the other equation...

myininaya (myininaya):

4x+5y=17 solve for y we subtract 4x on both sides 5y=-4x+17 now divide by 5 y=(-4x+17)/5 y=-4x/5+17/5

myininaya (myininaya):

to see where the lines intersect set them equal after putting both in y=mx+b form so we have -x/5-7/5=-4x/5+17/5 we can get rid of the fractions by multiplying both sides by 5 so we have -x-7=-4x+17

OpenStudy (anonymous):

this one seems more complicated.

myininaya (myininaya):

solve the above this equation for x by first add 4x on both sides 3x-7=17 adding 7 3x=24 x=8 to find out what y is we just plug into either y=-4x/5+17/5 or y=-x/5-7/5 so if x=8, y=-8/5-7/5=-15/5=-3 so the ordered pair is (x,y)=(8,-3)

OpenStudy (anonymous):

ahhh. but i like the other way better. this one seems like more work. but i find the other one easier.

OpenStudy (anonymous):

wait, can you see if im right on this other one i did.

OpenStudy (anonymous):

y= 1/2x - 2 ( a half times x minus 2 ) y= -3x +5

OpenStudy (anonymous):

i got (1, -2 )

OpenStudy (anonymous):

i plotted these equations and choose the ordered pair below where these lines intersect.

myininaya (myininaya):

x/2-2=-3x+5 x-4=-6x+10 7x=14 x=2 so y=-3(2)+5=-1 (x,y)=(2,-1)

OpenStudy (anonymous):

where did x/2 come from?

myininaya (myininaya):

isn't that 1/2 * x up there?

OpenStudy (anonymous):

nevermind. i got it.

OpenStudy (anonymous):

can you explain to me how you got x-4 = -6x + 10 what did you multiply?

myininaya (myininaya):

i just got rid of the fractions by multiplying both sides by 2

OpenStudy (anonymous):

hmmm. so you multiplied the reciprocal of x/2

OpenStudy (anonymous):

which is 2/x then you multiplied that by -2 and -3 and positive 5.

OpenStudy (anonymous):

and thats how you got your answer right?

myininaya (myininaya):

no i multiplied by 2 not 2/x

myininaya (myininaya):

2(x/2)-2(2)=2(-3x)+2(5) x-4=-6x+10

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

thank you so much!

myininaya (myininaya):

np

OpenStudy (anonymous):

can you please help me with another one?

OpenStudy (anonymous):

im not good in algebra 1 im only in 8th grade and this is hs stuff.

OpenStudy (anonymous):

hello.